OpenStudy (anonymous):

Find 4 consecutive odd integers with a sum of 184.

OpenStudy (anonymous):

\[n+n+2+n+4+n+6=184\] \[4n+12=184\] then done right?

OpenStudy (anonymous):

isnt that even integers?

OpenStudy (anonymous):

184-12=172/4=43

OpenStudy (anonymous):

43+45+47+49

OpenStudy (anonymous):

43+43+43+43 does not equal 184

OpenStudy (anonymous):

x- 1st number (x)+(x+2)+(x+4)+(x+6)=184 4x+12=184 4x=172 x=43 the nos are 43,45,47,49

OpenStudy (anonymous):

THanks very much

OpenStudy (anonymous):

ur welcome